# Riccardo Giordana Pozzi

### v1

@@ 3.3 Obtaining the full NLO contributions
That said, the results of such analysis can still be incorporated into the previous discussion. In particular, if we manage to determine the correct normalization in a single case for which universality applies, $\hat{\epsilon}_{\mathcal{T}}\hat{f}^{(1)}(z)$, we immediately identify the correct combination that can be applied consistently in all other situations.
Consider the scenario where considering explicitly a theory and we managed to fix $\hat{\epsilon}_{\mathcal{T}}$, by one of the above methods, to be $\hat{\epsilon}_{\mathcal{T}}=\frac{1}{4\pi^{2}C_{\mathcal{T}}}$. This will be the actual one, see subsection 4.1. The expression for the superprimary four-point function would be
@@
Notice also that the parameter of the expansion must be, for consistence, the same for all the superdisplacement components. This implies that if we consider a four-point function of the $Q$-descendants in the supermultiplet, these are related by supersymmetry to ${\hat{f}_{\mathbb{P}^{4}}^{(0)}}(z)$ and $\hat{f}_{\mathbb{P}^{4}}^{(1)}(z)$. Thus one can write explicitly the four-point function as given by
@@
What we may notice is that the $\epsilon_{\mathcal{T}}$ depends only on $C_{\mathcal{T}}$ and not $\mathbb{n}_{\mathbb{D}}$. This is true also for all the other four-point functions in the multiplet. So, $\epsilon_{\mathcal{T}}(C_{\mathcal{T}})$ is universal too and can be identified together with $\hat{f}_{\mathbb{D}^{4}}^{(1)}(z)$.
This implies, for example, that if the theory $\mathcal{T}$} is such that satisfy all the three conditions, if we have another theory having another line defect, say $\mathcal{T}$$^{\prime}$, also satisfying all the conditions, its displacement four-point function must be specified by
@@
When the defects have the same codimension this is identification is typically straightforward, when instead the codimension of the two line defects, $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$, is different the equivalence is true upon decomposing the tensor structures, see subsection 4.2.
The same holds also for all the other defect four-point function when ${\mathcal{D}}\notin{\mathcal{D}}\times{\mathcal{D}}$ and can be generalized for other cases, as discussed before, when relaxing either the second or the third condition.
It is worth adding a concluding remark regarding the holographic interpretation of dual theory, when available. Recall that we fixed the product $\epsilon_{\mathcal{T}}f^{(1)}(z)$ through explicit evaluation. However, due to the freedom of rescaling it may well be that the true parameter matching the holographic analysis is not $\epsilon_{\mathcal{T}}(C_{\mathcal{T}})$ itself, but rather $a\,\epsilon_{\mathcal{T}}(C_{\mathcal{T}})$ up to a rescaling by some constant $a$. What is guaranteed to match in explicit computations is not the parameter alone but the product of the two contributions $\hat{\epsilon}_{\mathcal{T}}\hat{f}^{(1)}(z)$
@@
4 Universal sectors in $1/2$ BPS Wilson lines
@@ 4.1 The $1/2$ BPS Wilson lines in ABJM and in ${\mathcal{N}}=4$ CSm theories
As a first example, we consider line defects of the same codimension. In particular, we take the $1/2$ BPS Wilson line in ABJM theory [21] and in ${\mathcal{N}}=4$ Chern–Simons–matter theories [22], under the framework discussed in section 3. The defect operators appearing in the two theories are characterized by the preserved superconformal algebra on the line. More precisely, they are classified by the quantum numbers of the bosonic subalgebra ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{su}}(3)_{{\mathcal{R}}}\oplus{\mathfrak{u}}(1)_{J_{0}}$ for the ABJM line, and ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{su}}(2)_{{\mathcal{R}}}\oplus{\mathfrak{u}}(1)_{J_{0}}$ for the SCSm one [34].
@@
$\begin{array}[]{lllll}{\mathcal{D}}\text{ in ABJM}&\qquad[{\textstyle\frac{3}{2}}]^{(0,0)}_{\frac{1}{2}}\longrightarrow&[2]^{(1,0)}_{1}\longrightarrow&\left[{\textstyle\frac{5}{2}}\right]_{\frac{3}{2}}^{(0,1)}\longrightarrow&[3]_{2}^{(0,0)}\;,\\ &&&&\\ {\mathcal{D}}\text{ in }{\mathcal{N}}=4\text{ SCSm}&\qquad&[1]_{1}^{(0)}\longrightarrow&\left[1\right]_{\frac{3}{2}}^{(1)}\longrightarrow&[1]_{2}^{(0)}\;,\end{array}$ (4.1)
where the charge assignment is schematically wirtten as $j_{0}]_{\Delta}^{(\mathcal{R})}$, with $(\mathcal{R})$ standing for the Dynkin labels of the considered representation and $j_{0}$ the abelian charge. For the two case the label for the R-symmetries refer to the ${\mathfrak{su}}(3)$ for the line in ABJM and ${\mathfrak{su}}(2)$ for the SCSm case. The arrows represent schematically the action of the supercharges in the respective theory.
It follows, from the discussion in section 3, that both theories satisfy the three assumptions we required. Hence, in both cases, when considering the four-point function of a single defect operator, no operators from the other component of the multiplets is exchanged. Consequently, the four-point are universal and hence the correlation function of one theory can be directly mapped to that of the other.
This observation is particularly useful when bootstrapping the four-point functions in SCSm. Indeed, imposing standard constraints such as crossing symmetry and consistency conditions does not suffice to completely fix the correlator, essentially due to the reduced supersymmetry that limits the problem. By contrast, once one notices the universal feature, one can bootstrap the full function up to NLO without resorting to the usual ansatz-driven bootstrap procedure.
For concreteness, let us consider the displacement four-point function (2.8). We focus on this case simply because the displacement operator is neutral under the R-symmetry in both theories, making the identification more transparent. The discussion, however, applies equally well to the other four-point functions, once one properly deals with the different representations, see for instance the next subsection.
The two displacement operators differ in their abelian charge, but this does not affect the $\mathbb{D}\times\bar{\mathbb{D}}$ expansion, which only selects neutral combinations. With this in mind, we can now look at the four-point function that reads
@@
for the line in ABJM. For the line in SCSm theories the four-point function will be, beside overall theory-specific constants, the very same expression
@@
In [22] the same conclusion was reached by observing that certain ${\mathcal{N}}=4$ CSm theories can be obtained from ABJM through specific orbifold constructions. Here we strengthen this result by showing that it indeed holds for all such $1/2$ BPS line in ${\mathfrak{osp}}(4|4)$ theories, that also satisfy the three requirements outlined in section 3.
@@
where now $C_{\text{SCSm}}$ is the normalization of the superfield and superprimary of the $1/2$ BPS defect of SCSm theories. For both of them we have, as anticipated
@@
In this case, we have preformed the identification for the displacement operator, but a symilar analysis applies also for all the other four-point function of defect operators in (4.1) that have the same conformal dimension.
@@ 4.2 The $1/2$ BPS Wilson line in ${\mathcal{N}}=4$ SYM
We now turn to another interesting case, namely the exploration of universality across defect lines of different codimension. In particular, we consider the $1/2$ BPS Wilson line in ABJM and in ${\mathcal{N}}=4$ SYM, building on the results obtained in [21, 8, 14, 19, 18]. The Wilson line in ${\mathcal{N}}=4$ SYM preserves the bosonic subalgebra ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{so}}(3)_{\text{rot}}\oplus\mathfrak{so}(5)_{\mathcal{R}}$ [^8] In [34] ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{su}}(2)_{\text{rot}}\oplus\mathfrak{usp}(4)_{\mathcal{R}}$. At the level of algebra this is equivalent given that, we recall, $\mathfrak{usp}(4)\simeq{\mathfrak{so}}(5)_{R}$ and ${\mathfrak{su}}(2)_{\text{rot}}\simeq{\mathfrak{so}}(3)_{\text{rot}}$.
In this case, the dual holographic description is well understood, and much of the discussion can indeed be supported by direct arguments from the bulk perspective. In this regard, one can perform very similar considerations to those presented in section 5. Here, however, we restrict to a purely CFT-based analysis.
@@
As previously introduced, the Dynkin labels in round and square brackets refer to the representations of ${\mathfrak{su}}(3)_{{\mathcal{R}}}$ and ${\mathfrak{u}}(1)_{j_{0}}$ for the Wilson line in ABJM, while they correspond to ${\mathfrak{so}}(5)_{R}$ and ${\mathfrak{so}}(3)_{\text{rot}}$ for the Wilson line in SYM [^9] See [34] for a detailed discussion of the representations.. Crucially, by studying the OPE relations for the superdisplacement multiplet [18, 19], one finds again that ${\mathcal{D}}\notin{\mathcal{D}}\times{\mathcal{D}}$, so all the required conditions are satisfied.
@@
such that $\mathbb{D}^{i}(t)$ transforms in the fundamental of $SO(2)$. The four-point function can thus be expressed as
$\langle\mathbb{D}(t_{1})\bar{\mathbb{D}}(t_{2})\mathbb{D}(t_{3})\bar{\mathbb{D}}(t_{4})\rangle_{\text{ABJM}}=\frac{1}{4}\sum\sigma_{ijkl}\langle\mathbb{D}^{i}(t_{1}){\mathbb{D}}^{j}(t_{2})\mathbb{D}^{k}(t_{3}){\mathbb{D}}^{l}(t_{4})\rangle_{\text{ABJM}}\,$, (4.10)
where $i,j,k,l=1,2$ and $\sigma_{ijkl}=\delta_{ij}\delta_{kl}+\delta_{il}\delta_{jk}-\delta_{ik}\delta_{jl}$, and the correlators on the right-hand side can be written as
$\langle\mathbb{D}^{i}(t_{1}){\mathbb{D}}^{j}(t_{2})\mathbb{D}^{k}(t_{3}){\mathbb{D}}^{l}(t_{4})\rangle_{\text{ABJM}}=\frac{{\mathcal{C}}_{\text{ABJM}}^{2}}{t_{12}^{4}t_{34}^{4}}F^{ijkl}(z)\,$, (4.11)
for some function $F^{ijkl}(z)$. For convenience, here we have absorbed all the overall factors in ${\mathcal{C}}_{\text{ABJM}}$.
Following the analysis in [8], it is convenient to re-express (4.11) in terms of the singlet $F_{S}(z)$, the symmetric traceless $F_{T}(z)$, and the antisymmetric $F_{A}(z)$ contributions associated with the two pairs of indices. This is done by taking the decomposition
$F^{ijkl}(z)=\delta_{ij}\delta_{kl}\,F_{S}(z)+(\delta_{ij}\delta_{kl}-\delta_{il}\delta_{jk})F_{A}(z)+(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\,\delta_{ij}\delta_{kl})F_{T}(z)\,.$ (4.12)
@@
$G^{ijkl}_{SO(3)}(z)=\delta_{ij}\delta_{kl}\,G_{S}(z)+(\delta_{ij}\delta_{kl}-\delta_{il}\delta_{jk})G_{A}(z)+(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-2/3\,\delta_{ij}\delta_{kl})G_{T}(z)\,.$ (4.14)
Crucially, in order to compare the two expression we need to decompose (4.14) in $SO(2)$ representation. This leads to the expression
$G^{ijkl}_{SO(2)}(z)=\delta_{ij}\delta_{kl}\,(G_{S}(z)+1/3G_{T}(z))+(\delta_{ij}\delta_{kl}-\delta_{il}\delta_{jk})G_{A}(z)+(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\,\delta_{ij}\delta_{kl})G_{T}(z)\,$, (4.15)
which is precisely on the form of (4.12), upon the identifications $F_{S}(z)=G_{S}(z)+1/3\,G_{T}(z)$, $F_{T}(z)=G_{T}(z)$ and $F_{A}(z)=G_{A}(z)$.
Based on the preliminary considerations, we expect universal features up to NLO, which implies a matching between the two four-point functions $F^{ijkl}(z)=G^{ijkl}_{SO(2)}(z)$.
@@
where we kept $\chi=\frac{z}{z-1}$ for keeping the expression concise [^10] Compared to [8, 14] we have rescaled these function by a factor of 2 to have the discussion more straightforward. See (4.24) .
@@
We end this subsection with the explicit check that, through this analysis, we obtain not only the NLO function but indeed the full NLO contribution, including the correct expansion parameter. From this perspective, let us consider the full expressions in the form of (3.17), where we explicitly take $\epsilon_{\mathcal{T}}$ as given in (4.8). These are
@@
Accordingly, equation (4.23) can be written explicitly as
@@
which, when evaluated explicitly, matches the expressions in [8, 14] which are written equivalently in terms of $\tilde{G}_{(1)}^{ijkl}(z)$, under the following redefinition
@@
As argued at the end of subsection 3, this reflects the fact that, through the identification of the universal sectors, we are matching the entire NLO contribution rather than isolating the expansion parameter $\epsilon_{\mathcal{T}}$ itself. The latter can always be rescaled by a constant, which can then be absorbed into the definition of the NLO function, as illustrated in the example above.

### v2

@@ 3.3 Obtaining the full NLO contributions
That said, the results of such an analysis can still be incorporated into the previous discussion. In particular, if we manage to determine the correct normalization in a single case for which universality applies, $\hat{\epsilon}_{\mathcal{T}}\hat{f}^{(1)}(z)$, we immediately identify the correct combination that can be applied consistently in all other situations.
Consider the scenario where considering explicitly a theory and we managed to fix $\hat{\epsilon}_{\mathcal{T}}$, by one of the above methods, to be $\hat{\epsilon}_{\mathcal{T}}=\frac{1}{4\pi^{2}C_{\mathcal{T}}}$. This will be the actual one; see Sec. 4.1. The expression for the superprimary four-point function would be
@@
Notice also that the parameter of the expansion must be, for consistency, the same for all the superdisplacement components. This implies that if we consider a four-point function of the $Q$-descendants in the supermultiplet, these are related by supersymmetry to ${\hat{f}_{\mathbb{P}^{4}}^{(0)}}(z)$ and $\hat{f}_{\mathbb{P}^{4}}^{(1)}(z)$. Thus one can write explicitly the four-point function as given by
@@
What we may notice is that $\epsilon_{\mathcal{T}}$ depends only on $C_{\mathcal{T}}$ and not $\mathbb{n}_{\mathbb{D}}$. This is true also for all the other four-point functions in the multiplet. So, $\epsilon_{\mathcal{T}}(C_{\mathcal{T}})$ is universal too and can be identified together with $\hat{f}_{\mathbb{D}^{4}}^{(1)}(z)$.
This implies, for example, that if the theory $\mathcal{T}$} is such that it satisfies all the three conditions, if we have another theory having another line defect, say $\mathcal{T}$$^{\prime}$, also satisfying all the conditions, its displacement four-point function must be specified by
@@
When the defects have the same codimension this identification is typically straightforward, when instead the codimension of the two line defects, $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$, is different, the equivalence is true upon decomposing the tensor structures; see Sec. 4.2.
The same holds also for all the other defect four-point functions when ${\mathcal{D}}\notin{\mathcal{D}}\times{\mathcal{D}}$ and can be generalized for other cases, as discussed before, when relaxing either the second or the third condition.
It is worth adding a concluding remark regarding the holographic interpretation of dual theory, when available. Recall that we fixed the product $\epsilon_{\mathcal{T}}f^{(1)}(z)$ through explicit evaluation. However, due to the freedom of rescaling it may well be that the true parameter matching the holographic analysis is not $\epsilon_{\mathcal{T}}(C_{\mathcal{T}})$ itself, but rather $a\,\epsilon_{\mathcal{T}}(C_{\mathcal{T}})$ up to a rescaling by some constant $a$. What is guaranteed to match in explicit computations is not the parameter alone but the product of the two contributions $\hat{\epsilon}_{\mathcal{T}}\hat{f}^{(1)}(z)$.
@@
4 Universal sectors in $1/2$ BPS line defects
@@ 4.1 The $1/2$ BPS Wilson lines in ABJM and in ${\mathcal{N}}=4$ CSm theories
As a first example, we consider line defects of the same codimension. In particular, we take the $1/2$ BPS Wilson line in ABJM theory [7] and in ${\mathcal{N}}=4$ Chern-Simons-matter theories [52], under the framework discussed in Sec. 3. The defect operators appearing in the two theories are characterized by the preserved superconformal algebra on the line. More precisely, they are classified by the quantum numbers of the bosonic subalgebra ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{su}}(3)_{{\mathcal{R}}}\oplus{\mathfrak{u}}(1)_{J_{0}}$ for the ABJM line, and ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{su}}(2)_{{\mathcal{R}}}\oplus{\mathfrak{u}}(1)_{J_{0}}$ for the ${\mathcal{N}}=4$ CSm one [2].
@@
$\begin{array}[]{lllll}{\mathcal{D}}\text{ in ABJM}&\qquad[{\textstyle\frac{3}{2}}]^{(0,0)}_{\frac{1}{2}}\longrightarrow&[2]^{(1,0)}_{1}\longrightarrow&\left[{\textstyle\frac{5}{2}}\right]_{\frac{3}{2}}^{(0,1)}\longrightarrow&[3]_{2}^{(0,0)}\;,\\ &&&&\\ {\mathcal{D}}\text{ in }{\mathcal{N}}=4\text{ CSm}&\qquad&[1]_{1}^{(0)}\longrightarrow&\left[1\right]_{\frac{3}{2}}^{(1)}\longrightarrow&[1]_{2}^{(0)}\;,\end{array}$ (4.1)
where the charge assignment is schematically written as $j_{0}]_{\Delta}^{(\mathcal{R})}$, with $(\mathcal{R})$ standing for the Dynkin labels of the considered representation and $j_{0}$ the Abelian charge. For the two cases the labels for the R-symmetries refer to the ${\mathfrak{su}}(3)$ for the line in ABJM and ${\mathfrak{su}}(2)$ for the ${\mathcal{N}}=4$ CSm case. The arrows represent schematically the action of the supercharges in the respective theories.
It follows, from the discussion in Sec. 3, that both theories satisfy the three assumptions we required. Hence, in both cases, when considering the four-point function of a single defect operator, no operators from the other component of the multiplets is exchanged. Consequently, the four-point functions are universal up to NLO and hence the correlation function of one theory can be directly mapped to that of the other.
This observation is particularly useful when bootstrapping the four-point functions in ${\mathcal{N}}=4$ CSm. Indeed, imposing standard constraints such as crossing symmetry and consistency conditions does not suffice to completely fix the correlator, essentially due to the reduced supersymmetry that limits the problem. By contrast, once one notices the universal feature, one can bootstrap the full function up to NLO without resorting to the usual ansatz-driven bootstrap procedure.
For concreteness, let us consider the displacement four-point function (2.9). We focus on this case because the displacement operator is neutral under the R-symmetry in both theories, which makes the identification particularly transparent. The discussion, however, applies more generally to correlation functions of compatible operators, namely operators with the same conformal dimension but transforming in different representations, e.g. the tilt operators. In such cases, one must first choose an appropriate embedding in order to place the correlators on the same footing as discussed in the next sections.
In the presence of chiral and antichiral superdisplacement multiplets, the displacement operators are naturally organized into complex combinations that make the $U(1)$ symmetry manifest. From this perspective, while the two displacement operators differ by their Abelian charge, this distinction plays no role in the $\mathbb{D}\times\bar{\mathbb{D}}$ OPE, which involves only neutral combinations. With this in mind, we can now look at the four-point function that reads
@@
for the line in ABJM. For the line in ${\mathcal{N}}=4$ CSm (SCSm) theories the four-point function will be, besides overall theory-specific constants, the very same expression
@@
In [52] the same conclusion was reached by observing that certain ${\mathcal{N}}=4$ CSm theories can be obtained from ABJM through specific orbifold constructions. Here we strengthen this result by showing that it indeed holds for all such $1/2$ BPS line in ${\mathfrak{osp}}(4|4)$ theories, that also satisfy the three requirements outlined in Sec. 3.
@@
where now $C_{\text{SCSm}}$ is the normalization of the superfield and superprimary of the $1/2$ BPS defect of ${\mathcal{N}}=4$ CSm theories. For both of them we have, as anticipated
@@
In this case, we have preformed the identification for the displacement operator, but a similar analysis applies also for all the other four-point functions of defect operators in (4.1) that have the same conformal dimension.
@@ 4.2 The $1/2$ BPS Wilson line in ${\mathcal{N}}=4$ SYM
We now turn to another interesting case, namely the exploration of universality across defect lines of different codimensions. In particular, we consider the $1/2$ BPS Wilson line in ABJM and in ${\mathcal{N}}=4$ SYM, building on the results obtained in [7, 41, 46, 35, 34]. The Wilson line in ${\mathcal{N}}=4$ SYM preserves the bosonic subalgebra ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{so}}(3)_{\text{rot}}\oplus\mathfrak{so}(5)_{\mathcal{R}}$. [^9] In [2] ${\mathfrak{so}}(2,1)_{\text{conf}}\oplus{\mathfrak{su}}(2)_{\text{rot}}\oplus\mathfrak{usp}(4)_{\mathcal{R}}$. At the level of algebra this is equivalent given that, we recall, $\mathfrak{usp}(4)\simeq{\mathfrak{so}}(5)_{R}$ and ${\mathfrak{su}}(2)_{\text{rot}}\simeq{\mathfrak{so}}(3)_{\text{rot}}$
In this case, the dual holographic description is well understood, and much of the discussion can indeed be supported by direct arguments from the bulk perspective. In this regard, one can perform very similar considerations to those presented in Sec. 5. Here, however, we restrict the discussion to a purely CFT-based analysis.
@@
As previously introduced, the Dynkin labels in round and square brackets refer to the representations of ${\mathfrak{su}}(3)_{{\mathcal{R}}}$ and ${\mathfrak{u}}(1)_{j_{0}}$ for the Wilson line in ABJM, while they correspond to ${\mathfrak{so}}(5)_{R}$ and ${\mathfrak{so}}(3)_{\text{rot}}$ for the Wilson line in SYM. [^10] See [2] for a detailed discussion of the representations. Crucially, by studying the OPE relations for the superdisplacement multiplet [34, 35], one finds again that ${\mathcal{D}}\notin{\mathcal{D}}\times{\mathcal{D}}$, so all the required conditions are satisfied.
@@
such that $\mathbb{D}^{\mathrm{i}}(t)$ transforms in the fundamental of $SO(2)$. The four-point function can thus be expressed as
$\langle\mathbb{D}(t_{1})\bar{\mathbb{D}}(t_{2})\mathbb{D}(t_{3})\bar{\mathbb{D}}(t_{4})\rangle_{\text{ABJM}}=\frac{1}{4}\sum\sigma_{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}\langle\mathbb{D}^{\text{i}}(t_{1}){\mathbb{D}}^{\mathrm{j}}(t_{2})\mathbb{D}^{\mathrm{k}}(t_{3}){\mathbb{D}}^{\mathrm{l}}(t_{4})\rangle_{\text{ABJM}}\,$, (4.10)
where $\mathrm{i},\mathrm{j},\mathrm{k},\mathrm{l}=1,2$ and $\sigma_{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}=\delta_{\mathrm{i}\mathrm{j}}\delta_{\mathrm{k}\mathrm{l}}+\delta_{\mathrm{i}\mathrm{l}}\delta_{\mathrm{j}\mathrm{k}}-\delta_{\mathrm{i}\mathrm{k}}\delta_{\mathrm{j}\mathrm{l}}$, and the correlators on the right-hand side can be written as
$\langle\mathbb{D}^{\text{i}}(t_{1}){\mathbb{D}}^{\mathrm{j}}(t_{2})\mathbb{D}^{\mathrm{k}}(t_{3}){\mathbb{D}}^{\mathrm{l}}(t_{4})\rangle_{\text{ABJM}}=\frac{{\mathcal{C}}_{\text{ABJM}}^{2}}{t_{12}^{4}t_{34}^{4}}F^{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}(z)\,$, (4.11)
for some function $F^{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}(z)$. For convenience, here we have absorbed all the overall factors in ${\mathcal{C}}_{\text{ABJM}}$.
Following the analysis in [41], it is convenient to reexpress (4.11) in terms of the singlet $F_{S}(z)$, the symmetric traceless $F_{T}(z)$, and the antisymmetric $F_{A}(z)$ contributions associated with the two pairs of indices. This is done by taking the decomposition
$F^{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}(z)=\delta^{\mathrm{i}\mathrm{j}}\delta^{\mathrm{k}\mathrm{l}}\,F_{S}(z)+(\delta^{\mathrm{i}\mathrm{j}}\delta^{\mathrm{k}\mathrm{l}}-\delta^{\mathrm{i}\mathrm{l}}\delta^{\mathrm{j}\mathrm{k}})F_{A}(z)+(\delta^{\mathrm{i}\mathrm{k}}\delta^{\mathrm{j}\mathrm{l}}+\delta^{\mathrm{i}\mathrm{l}}\delta^{\mathrm{j}\mathrm{k}}-\,\delta^{\mathrm{i}\mathrm{j}}\delta^{\mathrm{k}\mathrm{l}})F_{T}(z)\,.$ (4.12)
@@
$G^{ijkl}_{SO(3)}(z)=\delta^{ij}\delta^{kl}\,G_{S}(z)+(\delta^{ij}\delta^{kl}-\delta^{il}\delta^{jk})G_{A}(z)+(\delta^{ik}\delta^{jl}+\delta^{il}\delta^{jk}-2/3\,\delta^{ij}\delta^{kl})G_{T}(z)\,.$ (4.14)
Crucially, in order to compare the two expressions we need to decompose (4.14) into $SO(2)$ representations. This can be done by using the standard embedding of $SO(2)$ into $SO(3)$, realized as rotations about the third axis. This leads to the expression
$G^{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}_{SO(2)}(z)=\delta^{\mathrm{i}\mathrm{j}}\delta^{\mathrm{k}\mathrm{l}}\,(G_{S}(z)+1/3G_{T}(z))+(\delta^{\mathrm{i}\mathrm{j}}\delta^{\mathrm{k}\mathrm{l}}-\delta^{\mathrm{i}\mathrm{l}}\delta^{\mathrm{j}\mathrm{k}})G_{A}(z)+(\delta^{\mathrm{i}\mathrm{k}}\delta^{\mathrm{j}\mathrm{l}}+\delta^{\mathrm{i}\mathrm{l}}\delta^{\mathrm{j}\mathrm{k}}-\,\delta^{\mathrm{i}\mathrm{j}}\delta^{\mathrm{k}\mathrm{l}})G_{T}(z)\,$, (4.15)
which is precisely on the form of (4.12), upon the identifications
$F_{S}(z)=G_{S}(z)+1/3\,G_{T}(z)\,,\qquad F_{T}(z)=G_{T}(z)\,,\quad\text{and}\quad F_{A}(z)=G_{A}(z)\,.$
Based on the preliminary considerations, we expect universal features up to NLO, which implies a matching between the two four-point functions $F^{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}(z)=G^{\mathrm{i}\mathrm{j}\mathrm{k}\mathrm{l}}_{SO(2)}(z)$.
@@
where we kept $\chi=\frac{z}{z-1}$ for keeping the expression concise. [^11] Compared to [41, 46] we have rescaled these function by a factor of 2 to have the discussion more straightforward. See (4.24)
@@
We end this section with the explicit check that, through this analysis, we obtain not only the NLO function but indeed also the full NLO contribution, including the correct expansion parameter. From this perspective, let us consider the full expressions in the form of (3.17), where we explicitly take $\epsilon_{\mathcal{T}}$ as given in (4.8). These are
@@
Accordingly, Eq. (4.23) can be written explicitly as
@@
which, when evaluated explicitly, matches the expressions in [41, 46] which are written equivalently in terms of $\tilde{G}_{(1)}^{ijkl}(z)$, under the following redefinition:
@@
As argued at the end of Sec. 3, this reflects the fact that, through the identification of the universal sectors, we are matching the entire NLO contribution rather than isolating the expansion parameter $\epsilon_{\mathcal{T}}$ itself. The latter can always be rescaled by a constant, which can then be absorbed into the definition of the NLO function, as illustrated in the example above.